# Nonlocal Locking of Observable Quantities: A Faithful Signature of Nonclassical Correlations

### v1

@@ Abstract
Nonclassicality in composite quantum systems depicts several puzzling manifestations, with Einstein-Podolsky-Rosen entanglement, Schrödinger steering, and Bell nonlocality being the most celebrated ones. In addition to those, an unentangled quantum state can also exhibit nonclassicality, as evidenced from notions such as quantum discord and work deficit. Here, we propose a general framework to investigate nonclassical correlations in multipartite quantum states. The distinct signatures left on observable quantities, depending on whether the sub-parts of a composite system are probed separately or jointly, provide an operational avenue to construct different quantifiers that faithfully capture signatures of nonclassicality in quantum states. Along the line we unveil an intriguing phenomenon referred to as ‘nonlocal locking of observable quantities’, where the value of an observable quantity gets locked in the correlation of a nonclassical state. Our approach reduces the experimental demand for verification of nonclassicality in composite systems and can find applications for enhaced energy storage in quantum thermodynamical devices.
@@
Introduction.– Correlations stand at the forefront of scientific inquiry across various fields of study. Their systematic investigation plays a crucial role in enhancing our understanding of the physical world [1]. This is especially evident while studying quantum systems consisting of multiple sub-parts. For instance, the puzzling features of quantum correlations as exemplified by the phenomena of entanglement and steering have been recognized since the early days of quantum theory [2, 3, 4, *Schrdinger1936]. Later J. S. Bell, in his seminal work, built on the notion of quantum entanglement to establish the physical phenomenon of quantum nonlocality [6, *Bell1966] (see also [8]). Along with capturing fundamentally distinct notions of nonclassicality [9, 10, 11], quantum entanglement, steering, and nonlocality also find a number of utilities across a spectrum of applications at the heart of modern quantum technologies [12, 13, 14, 15]. More recently, it has been identified that quantum discord encapsulates nonclassicality even in an unentangled quantum state [16, 17, 18] (see also [19, 20]). The study of nonclassical correlations, with an emphasis on their identification, quantification, and classification, thus stands as a pivotal objective in quantum information research during the last few decades.
In this paper, we adopt an operationally-driven approach to explore nonclassical correlations in multipartite quantum systems. A multipartite quantum system can be probed in different ways depending on the availability of resources among the spatially separated sub-parts. For instance, global probing requires quantum communication channels among the sub-parts, and without such channels only local probing is possible, where distant parties address their respective sub-parts individually. On the other hand, allowing only classical communication among the sub-parts confines a probing mechanism intermediate to the global and local probings (see Fig. 1). A correlated quantum state can lead to distinct observable values depending on how the system is probed and this observation provides operational pathways for exploring nonclassical correlations within it. In this work, we pursue this approach to introduce quantifiers of general nonclassical correlations beyond entanglement. While some quantifiers are derived from resource theories and hinge on entropic quantities, the generality of our framework allows for constructing alternative quantifiers leveraging simple observables on composite systems, thus mitigating experimental complexity. For bipartite systems, such quantifiers are proven to be faithful, meaning that they vanish on and only on states lacking nonclassical correlations. Our analysis unveils an intriguing phenomenon hereby termed ‘nonlocal locking of observable quantities,’ wherein observable values get locked within the nonclassical correlation inherent in multipartite quantum states. Furthermore, we provide exact quantitative results focusing on the simplest composite quantum system comprising two qubits. Finally, we discuss on the potential implications of our findings in designing efficient energy storage devices that utilize nonclassically correlated states of many-body quantum systems.
@@
Preliminaries.– The state of an individual quantum system ($A$) is described by a density operator $\rho\in\mathcal{D}(\mathcal{H}_{A})$; $\mathcal{D}(\mathcal{X})$ be the set of positive trace-one operators acting on the Hilbert space $\mathcal{X}$. For a finite-dimensional system, which we assume throughout this work, $\mathcal{H}_{A}$ is isomorphic to some complex Euclidean space $\mathbb{C}^{d_{A}}$, where $d_{A}$ denotes the dimension of the Hilbert space. The state of a composite system ${\bf A}\equiv A_{1}\cdots A_{n}$ is called separable if it is a statistical mixture of independently prepared constituent sub-parts, i.e., $\rho_{\bf A}=\sum_{i}p_{i}(\otimes_{j=1}^{n}\psi^{i}_{j})$, where $\vec{p}:=\{p_{i}\}$ is a probability vector, and $\psi:=\ket{\psi}\bra{\psi}$ | with |$\ket{\psi^{i}}_{j}\in\mathbb{C}^{d_{j}}$. States that are not separable are called entangled, and they capture the most celebrated aspect of nonclassical correlation in composite quantum systems [12, 13]. Importantly, a separable state can also accommodate another distinct nonclassical aspect of correlation recognized as quantum discord [16, 17, 18]. A multipartite state is called classically correlated if it allows a decomposition of the form $\rho_{\bf A}=\sum p_{i_{1}\cdots i_{n}}\psi^{i_{1}}_{1}\otimes\cdots\otimes\psi^{i_{n}}_{n}$, where {|$\{\ket{\psi^{i_{j}}}_{j}\}$ forms an orthonormal basis (ONB) for the $j^{th}$ Hilbert space, and $\{p_{i_{1}\cdots i_{n}}\}$ denotes a probability vector. A function $\mathbb{F}:\mathcal{D}(\otimes_{j=1}^{n}\mathbb{C}^{d_{j}}_{A_{j}})\mapsto\mathbb{R}$ is said to capture the signature of nonclassical correlation faithfully if it takes value zero for all classically correlated states and nonzero otherwise. For the bipartite case, the original entropic measure of quantum discord, defined in terms of the difference between two different expressions for the mutual information [16, 17, 18], is an example of a faithful quantifier [20]. The notion of quantum discord has been generalized to quantum systems involving more than two sub-parts, albeit with caveats in its operational interpretation [21, 22, 23]. Notably, this notion of nonclassicality is party asymmetric, as a state can be a statistical mixture of product states where {|$\{\ket{\psi^{i_{j}}}\}$ form ONBs only for some sub-parts, but not for all. In the following, we develop a general operational framework for studying nonclassical correlations in multipartite quantum states.
@@
Framework.– Any physical operation acting on a quantum system is described by a completely positive trace-preserving (CPTP) map [24]. Given a quantum system $A$, we start by defining a generic function $f_{A}:\mathcal{D}(\mathbb{C}^{d_{A}})\mapsto\mathbb{R}$. For a CPTP map $\phi_{A}:\mathcal{D}(\mathbb{C}^{d_{A}})\mapsto\mathcal{D}(\mathbb{C}^{d_{A}})$, we define the variation of the function $f_{A}$ on some state $\rho_{A}$ as $\Delta f_{A}^{\phi_{A}}(\rho_{A}):=f_{A}(\rho_{A})-f_{A}(\phi(\rho_{A}))$. For a set of CPTP maps $\Phi_{A}$, the variation is given by $\Delta f_{A}^{\Phi_{A}}(\rho_{A}):=\max_{\phi_{A}\in\Phi_{A}}\Delta f_{A}^{\phi_{A}}(\rho_{A})$, which quantifies the maximum extracted value of the function $f_{A}$ on the given state $\rho_{A}$ under the set of operations $\Phi_{A}$. For a multipartite quantum system ${\bf A}(=A_{1}\cdots A_{n})$, fixing $(f_{j},\Phi_{j})$ pair for the $j^{th}$ party $j\in\{1,\cdots,n\}\equiv[n]$, and fixing $(f_{\bf A},\Phi_{\bf A})$ pair for the joint system, we define another function $\mathbb{F}_{GL}(\rho_{\bf A}):=\Delta f^{\Phi_{\bf A}}_{\bf A}(\rho_{\bf A})-\sum_{j}\Delta f^{\Phi_{j}}_{j}(\rho_{j})$, where $\rho_{j}:=\operatorname{Tr}_{A_{j}^{c}}(\rho_{\bf A})$ denotes the marginal state of the $j^{th}$ sub-part. The function $\mathbb{F}_{GL}$ quantifies the difference between the extracted value of the global function and that of the local functions evaluated on the global state and the local sub-parts under the fixed sets of operations. Interestingly, as we will see, for suitable choices of $(f_{\gamma},\Phi_{\gamma})$’s, with $\gamma\in\{[n]\cup{\bf A}\}$, the function $\mathbb{F}_{GL}$ can capture the signature of different forms of correlation within the multipartite state. As an example, for bipartite systems, the following observation is noteworthy (see Appendix-I for more details).
@@
Resource theoretic quantifiers.– To see application of the aforesaid general framework we start by analyzing how the function $\mathbb{F}_{G\mathcal{N}}$ can capture the notion of quantum discord in bipartite systems (${\bf A}=A_{1}A_{2}$) for suitable choices of the tuple $(f_{\gamma},\Phi_{\gamma},\mathcal{N})$. For that once again we fix $(f_{\gamma},\Phi_{\gamma})\equiv(f^{pe}_{\gamma},\Phi_{\gamma}^{no}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \gamma\in\{1,2,{\bf A}\}$, and fix $\mathcal{N}$ to be classical communication channel between $A_{1}$ and $A_{2}$. Recall that quantum discord is a party asymmetric concept: a bipartite state $\rho_{\bf A}\in\mathcal{D}(\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}})$ having the classical-quantum (CQ) form $\rho_{\bf A}=\sum p_{i_{1}i_{2}}\psi^{i_{1}}_{1}\otimes\psi^{i_{2}}_{2}$, with $\langle\psi^{i_{1}}_{1}|\psi^{i^{\prime}_{1}}_{1}\rangle=\delta_{i_{i}i^{\prime}_{1}}$, has discord zero from $A_{1}\to A_{2}$ [16, 17, 18]. A generic classical communication channel $N\in\mathcal{N}_{C}$ sending the $A_{1}$ part to the $A_{2}$ can be thought as a measure-and-prepare channel [30, 31], with preparations being restricted to rank-1 orthogonal states (see Appendix III). Such a channel could map completely mixed state to a resourceful state, and therefore is costly from the perspective of purity resource theory. One can bypass this cost association by further limiting these channels to be cost free. As we argue in Appendix III, a cost-free classical channel in this case takes the form $N(\rho_{A_{1}}):=\sum_{l\in[d_{1}]}\operatorname{Tr}(\pi_{l}\rho_{1})\ket{\alpha^{l}_{1}}\bra{\alpha^{l}_{1}}$ |, where $\langle\alpha^{l}_{1}|\alpha^{l^{\prime}}_{1}\rangle=\delta_{ll^{\prime}}$, and $\{\pi_{l}\}$ is a positive-operator-value-measurement (POVM) with $\operatorname{Tr}(\pi_{l})=1,\leavevmode\nobreak\ \forall\leavevmode\nobreak\ l\in[d_{1}]$. By denoting the set of all such zero cost classical channels as $\mathcal{N}^{0}_{C}$, we obtain the following result.
@@
Generic nonclassicality quantifiers.– We now proceed to explore the generality of our framework to construct quantifiers for nonclassical correlations that need not to be motivated from a resource theory. To this aim, we start by noting that the function $f:\mathcal{D}(\mathbb{C}^{d})\mapsto\mathbb{R}$ can simply be the expectation value of some observable $\mathcal{O}\in\mbox{Herm}(\mathbb{C}^{d})$, i.e., $\mathcal{O}:\rho\mapsto\langle\mathcal{O}\rangle_{\rho}:=\operatorname{Tr}(\mathcal{O}\rho)$. Unlike the earlier case, here the variation is evaluated over the set of reversible operations, resulting $\Phi$ to be the set of unitary operations $\mathrm{U}$. In this case the variation $\Delta\mathcal{O}^{\mathrm{U}}(\rho):=\max_{u\in\mathrm{U}}[\operatorname{Tr}(\mathcal{O}\rho)-\operatorname{Tr}(\mathcal{O}u\rho u^{\dagger})]$ quantifies the optimal value of the observable $\mathcal{O}$ extracted from the state $\rho$ when the system is evolved under closed system dynamics. While studying multipartite systems, generality of our framework allows choosing different observables for different sub-parts. Although the choice of the global observable can also be arbitrary, here we consider $\mathcal{O}_{\bf A}\equiv\sum_{j\in[n]}\tilde{\mathcal{O}}_{j}\leavevmode\nobreak$, where $\tilde{\mathcal{O}}_{j}:=\mathcal{O}_{j}\otimes\mathbf{I}_{rest}$. If system’s Hamiltonian $h$ is taken to be the observable, the variation $\Delta h^{\mathrm{U}}(\rho)$, known as the ergotropy of the state, which has a long history in literature [37, 38, 39]. The corresponding function $\mathbb{H}_{GL}(\rho_{\bf A})$ quantifies the ergotropic gap – the difference between the extractable work from a state $\rho_{\bf A}$ under global and local unitary operations, respectively. Interestingly, this function and its variants can capture different forms of nonclassical correlations, particularly entanglement in bipartite states and genuine entanglement in multipartite quantum states [40, 41, 42, 43, 44].
Figure 2: Nonlocal energy locking in two qubit Werner states $\mathcal{W}^{\alpha}_{\bf A}$. Red dashed curve denotes the original quantum discord $\mathcal{Q}(\mathcal{W}^{\alpha}_{\bf A})=\frac{1-\alpha}{4}\log(1-\alpha)-\frac{1+\alpha}{2}\log(1+\alpha)+\frac{1+3\alpha}{4}\log(1+3\alpha)$ establishing the states are nonclassical for $\alpha\in(0,1]$ [17] (see also [45]). Blue curve amounts to energy $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\mathcal{W}^{\alpha}_{\bf A})\leavevmode\nobreak\ [=\alpha\epsilon_{2}/2]$ locked into the nonclassical correlation of the state $\mathcal{W}^{\alpha}_{\bf A}$. Here the plot is given by fixing $\epsilon_{2}=2$.
@@
While studying the quantity $\mathbb{O}_{G\mathcal{N}^{0}_{C}}$, which is one of our core interests in the present work, we first need to identify the free classical channels $\mathcal{N}^{0}_{C}$. As it can be argued a classical channel $N\in\mathcal{N}_{C}$ communicating $A_{j}$ sub-part to the $j^{\star}$ party will be considered to be free if $\operatorname{Tr}(\mathcal{O}_{j}N(\rho_{j}))\leq\operatorname{Tr}(\mathcal{O}_{j}\rho_{j})\leavevmode\nobreak\ \&\leavevmode\nobreak\ \Delta\mathcal{O}^{U_{j}}_{j}N(\rho_{j})\leq\Delta\mathcal{O}^{U_{j}}_{j}(\rho_{j}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \rho_{j}\in\mathcal{D}(\mathbb{C}^{d_{j}})$. A more elaborate justification of this demand as well as a complete characterization of free classical channels $\mathcal{N}^{0}_{C}$ are provided in Appendix IV. With this we now present a general result for bipartite systems (proof is deferred to Appendix IV).
@@
The function $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ faithfully captures nonclassical correlation from $A_{1}\to A_{2}$ in any bipartite state $\rho_{\bf A}$, whenever $\mathcal{O}_{\bf A}$ has non degenerate spectrum.
Notably unlike Theorem 1, the function $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ faithfully captures signature of nonclassical correlation in any bipartite state without assuming any further constraint on the state. Theorem 2 has quite interesting implication. A non zero value of $\mathbb{O}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ implies that extractable value of the observable $\mathcal{O}_{\bf A}$ under global operations is strictly greater than the value extractable under local operations assisted with classical communication from $A_{1}$ to $A_{2}$. In other words, the value of the observable $\mathcal{O}_{\bf A}$ gets locked in the nonclassical correlation of the state. This phenomenon we call the nonlocal locking of observable quantities in nonclassical correlation. The motivation of this name arises from the pioneering ‘nonlocality without entanglement’ phenomenon, where distinguishability of a set of multipartite product states under local operation and classical communication is suboptimal than the global distinguishability [46, 47] (see also [48, 49, 50, 51]), and the works [52, 53, 54] that report several locking phenomena utilizing quantum resources). The following crucial property (proof provided in Appendix V) of the quantity discussed in Theorem 2 is worth noting.

### v3

# Ergodiscord: An Operational and Distinct Notion of Quantumness of Correlations

@@ Abstract
Nonclassicality in composite quantum systems depicts several puzzling manifestations, with Einstein-Podolsky-Rosen entanglement, Schrödinger steering, and Bell nonlocality being the most celebrated ones. In addition to those, an unentangled quantum state can also exhibit nonclassicality, as evidenced from notions such as quantum discord and work deficit. In this work, we present a general framework for exploring quantumness of correlations in multipartite quantum states. By exploiting the different signatures reflected on observable quantities depending on whether subsystems of a composite systems are probed jointly or independently, we introduce an operational quantifier of nonclassicality, termed ergodiscord. As we show, this newly proposed quantifier faithfully captures nonclassicality in any bipartite quantum state, while being fundamentally distinct from the original quantum discord. Moreover, ergodiscord uncovers an intriguing phenomenon called ‘nonlocal energy locking’, where a useful form of energy (i.e. work) gets locked in correlations of nonclassical states. We also show that a mixed nonclassical state can lock more work than the maximally entangled state of the corresponding system, establishing an interesting super-additivity phenomenon of nonlocal energy locking. The present work may inspire novel designs of quantum energy storage devices by utilizing nonclassical correlations in composite quantum systems.
I Introduction
@@
Correlations are central to scientific inquiry across many disciplines, with their systematic study playing a key role in advancing our understanding of the physical world [1]. This is especially evident in the study of quantum systems comprising multiple subsystems. The intriguing nature of quantum correlations, as exemplified by entanglement and steering, has been recognized since the early foundations of quantum theory [2, 3, 4, *Schrdinger1936]. Building on these ideas, J. S. Bell, in his landmark work, extended the concept of quantum entanglement to establish quantum nonlocality [6, *Bell1966] (see also [8]). Alongside their role in defining fundamental notions of nonclassicality [9, 10, 11], quantum entanglement, steering, and nonlocality have found numerous applications in the rapidly advancing field of quantum technologies [12, 13, 14, 15]. More recently, it has been shown that quantum discord captures nonclassicality even in unentangled states [16, 17, 18] (see also [19, 20]). Consequently, the investigation of nonclassical correlations, with a focus on their identification, quantification, and classification, has become a central objective in quantum information research over the past few decades.
In this paper, we adopt an operational approach to investigate nonclassical correlations in multipartite quantum systems. We start by noting that such systems can be probed differently depending on the resources available among the spatially separated subsystems. Global probing requires quantum communication channels among the different sub-parts, whereas in their absence, only local probing is feasible, where each subsystem is addressed individually. Allowing only classical communication enables an intermediate probing regime between the global and local cases (see Fig. 1). The distinct observable outcomes resulting from these different probing strategies provide a foundation for defining and quantifying nonclassical correlations. This operational framework offers a versatile methodology to introduce various quantifiers of nonclassicality. While this approach can be built on resource theories leading to entropic measures, our framework also enables the construction of alternative quantifiers using expectation values of observable quantities. In particular, we propose a novel thermodynamic-based quantity, called ‘ergodiscord’, which involves the works extractable from a multipartite system under reversible processes, that in general can be further constrained due to the availability of resources to implement them. As we show, for bipartite systems ergodiscord faithfully characterizes nonclassical correlations, vanishing only for states entirely devoid of such correlations. Our analysis also uncovers an operational phenomenon termed ‘nonlocal energy locking’, wherein extractable work from an isolated multipartite quantum system becomes locked within its nonclassical correlations. We derive exact expressions for ergodiscord in generic two-qubit states, demonstrating its distinctness from quantum discord [16, 17, 18] and suggesting that a resource theory of ergodiscord would differ qualitatively from that of quantum discord [19, 20]. Finally, we discuss how this study can contribute to the design of efficient energy storage devices by exploiting the quantumness of correlations in multipartite states, with potential applications in quantum batteries [21, 22, 23].
II Preliminaries
@@
The state of an individual quantum system ($A$) is described by a density operator $\rho\in\mathcal{D}(\mathcal{H}_{A})$. For a finite-dimensional system, which we will assume throughout this work, $\mathcal{H}_{A}$ is isomorphic to some complex Euclidean space $\mathbb{C}^{d_{A}}$, where $d_{A}$ denotes the dimension of the Hilbert space. The state of a composite system ${\bf A}\equiv A_{1}\cdots A_{n}$ is called separable if it is a statistical mixture of independently prepared constituent sub-parts, i.e., $\rho_{\bf A}=\sum_{i}p_{i}(\otimes_{j=1}^{n}\psi^{i}_{j})$, where $\vec{p}:=\{p_{i}\}$ is a probability vector, and $\psi:=\ket{\psi}\bra{\psi}$ | with |$\ket{\psi^{i}}_{j}\in\mathbb{C}^{d_{j}}$. States that are not separable are called entangled, and they capture the most celebrated aspect of nonclassical correlation in composite quantum systems [12, 13]. Importantly, a separable state can also accommodate nonclassical aspect recognized as quantum discord [16, 17, 18]. A multipartite state is called classically correlated if it allows a decomposition of the form $\rho_{\bf A}=\sum p_{i_{1}\cdots i_{n}}\psi^{i_{1}}_{1}\otimes\cdots\otimes\psi^{i_{n}}_{n}$, where {|$\{\ket{\psi^{i_{j}}}_{j}\}$ forms an orthonormal basis (ONB) for the $j^{th}$ Hilbert space, and $\{p_{i_{1}\cdots i_{n}}\}$ denotes a probability vector. A function $\mathbb{F}:\mathcal{D}(\otimes_{j=1}^{n}\mathbb{C}^{d_{j}}_{A_{j}})\mapsto\mathbb{R}$ is said to capture the signature of nonclassical correlation faithfully if it takes value zero for all classically correlated states and nonzero otherwise. For the bipartite case, the entropic measure of quantum discord, defined in terms of the difference between two different expressions for the mutual information [16, 17, 18], is an example of a faithful quantifier of nonclassical correlation [20]. Later, this notion has been generalized to quantum systems involving more than two sub-parts, albeit with caveats in its operational interpretation [24, 25, 26]. Notably, this notion of nonclassicality is party asymmetric, as a state can be a statistical mixture of product states, where {|$\{\ket{\psi^{i_{j}}}\}$ form ONBs only for some sub-parts, but not for all. In the following, we develop a general operational framework for studying nonclassical correlations in multipartite quantum states.
III Framework
@@
Any physical operation acting on a quantum system is described by a completely positive trace-preserving (CPTP) map [27]. Given a quantum system $A$, we start by defining a generic function $f_{A}:\mathcal{D}(\mathbb{C}^{d_{A}})\mapsto\mathbb{R}$. For a CPTP map $\phi_{A}:\mathcal{D}(\mathbb{C}^{d_{A}})\mapsto\mathcal{D}(\mathbb{C}^{d_{A}})$, we define the variation of the function $f_{A}$ on some state $\rho_{A}$ as $\Delta f_{A}^{\phi_{A}}(\rho_{A}):=f_{A}(\rho_{A})-f_{A}(\phi(\rho_{A}))$. For a set of CPTP maps $\Phi_{A}$, the variation is given by $\Delta f_{A}^{\Phi_{A}}(\rho_{A}):=\max_{\phi_{A}\in\Phi_{A}}\Delta f_{A}^{\phi_{A}}(\rho_{A})$, which quantifies the maximum extracted value of the function $f_{A}$ on the given state $\rho_{A}$ under the set of operations $\Phi_{A}$. For a multipartite quantum system ${\bf A}(=A_{1}\cdots A_{n})$, fixing $(f_{j},\Phi_{j})$ pair for the $j^{th}$ party $j\in\{1,\cdots,n\}\equiv[n]$, and fixing $(f_{\bf A},\Phi_{\bf A})$ pair for the joint system, we define another function $\mathbb{F}_{GL}(\rho_{\bf A}):=\Delta f^{\Phi_{\bf A}}_{\bf A}(\rho_{\bf A})-\sum_{j}\Delta f^{\Phi_{j}}_{j}(\rho_{j})$, where $\rho_{j}:=\operatorname{Tr}_{A_{j}^{c}}(\rho_{\bf A})$ denotes the marginal state of the $j^{th}$ sub-part. The function $\mathbb{F}_{GL}$ quantifies the difference between the extracted value of the global function and that of the local functions evaluated on the global state and the local sub-parts under the fixed sets of operations. Interestingly, as we will see, for suitable choices of $(f_{\gamma},\Phi_{\gamma})$’s, with $\gamma\in\{[n]\cup{\bf A}\}$, the function $\mathbb{F}_{GL}$ can capture the signature of different forms of correlation within the multipartite state. As an example, for bipartite systems, the following observation is noteworthy (see Appendix I for more details).
@@
IV Resource theoretic quantifiers
To see application of the aforesaid general framework we start by analyzing how the function $\mathbb{F}_{G\mathcal{N}}$ can capture the notion of quantum discord in bipartite systems (${\bf A}=A_{1}A_{2}$) for suitable choices of the tuple $(f_{\gamma},\Phi_{\gamma},\mathcal{N})$. For that once again we fix $(f_{\gamma},\Phi_{\gamma})\equiv(f^{pe}_{\gamma},\Phi_{\gamma}^{no}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \gamma\in\{1,2,{\bf A}\}$, and fix $\mathcal{N}$ to be classical communication channel between $A_{1}$ and $A_{2}$. Recall that quantum discord is a party asymmetric concept: a bipartite state $\rho_{\bf A}\in\mathcal{D}(\mathbb{C}^{d_{1}}\otimes\mathbb{C}^{d_{2}})$ having the classical-quantum (CQ) form $\rho_{\bf A}=\sum p_{i_{1}i_{2}}\psi^{i_{1}}_{1}\otimes\psi^{i_{2}}_{2}$, with $\langle\psi^{i_{1}}_{1}|\psi^{i^{\prime}_{1}}_{1}\rangle=\delta_{i_{i}i^{\prime}_{1}}$, has discord zero from $A_{1}\to A_{2}$ [16, 17, 18]. A generic classical communication channel $N\in\mathcal{N}_{C}$ sending the $A_{1}$ part to the $A_{2}$ can be thought as a measure-and-prepare channel [33, 34], with preparations being restricted to rank-1 orthogonal states (see Appendix III). Such a channel could map completely mixed state to a resourceful state, and therefore is costly from the perspective of purity resource theory. One can bypass this cost association by further limiting these channels to be cost free. As we argue in Appendix III, a cost-free classical channel in this case takes the form $N(\rho_{A_{1}}):=\sum_{l\in[d_{1}]}\operatorname{Tr}(\pi_{l}\rho_{1})\ket{\alpha^{l}_{1}}\bra{\alpha^{l}_{1}}$ |, where $\langle\alpha^{l}_{1}|\alpha^{l^{\prime}}_{1}\rangle=\delta_{ll^{\prime}}$, and $\{\pi_{l}\}$ is a positive-operator-value-measurement (POVM) with $\operatorname{Tr}(\pi_{l})=1,\leavevmode\nobreak\ \forall\leavevmode\nobreak\ l\in[d_{1}]$. By denoting the set of all such zero cost classical channels as $\mathcal{N}^{0}_{C}$, we obtain the following result.
@@ V Generic nonclassicality quantifiers and Ergodiscord
We now proceed to explore the generality of our framework to construct quantifiers for nonclassical correlations that need not to be motivated from a resource theory. To this aim, we start by noting that the function $f:\mathcal{D}(\mathbb{C}^{d})\mapsto\mathbb{R}$ can simply be the expectation value of some observable $\mathcal{O}\in\mbox{Herm}(\mathbb{C}^{d})$, i.e., $\mathcal{O}:\rho\mapsto\langle\mathcal{O}\rangle_{\rho}:=\operatorname{Tr}(\mathcal{O}\rho)$. Unlike the earlier case, here the variation is evaluated over the set of reversible operations, resulting $\Phi$ to be the set of unitary operations $\mathrm{U}$. In this case the variation $\Delta\mathcal{O}^{\mathrm{U}}(\rho):=\max_{u\in\mathrm{U}}[\operatorname{Tr}(\mathcal{O}\rho)-\operatorname{Tr}(\mathcal{O}u\rho u^{\dagger})]$ quantifies the optimal value of the observable $\mathcal{O}$ extracted from the state $\rho$ when the system is evolved under closed system dynamics. While studying multipartite systems, generality of our framework allows choosing different observables for different sub-parts. Although the choice of the global observable can also be arbitrary, here we consider $\mathcal{O}_{\bf A}\equiv\sum_{j\in[n]}\tilde{\mathcal{O}}_{j}\leavevmode\nobreak$, where $\tilde{\mathcal{O}}_{j}:=\mathcal{O}_{j}\otimes\mathbf{I}_{rest}$. From here onward, we, without loss of any generality, will consider the observable to be the Hamiltonian of the system. Accordingly, the variation $\Delta h^{\mathrm{U}}(\rho)$ becomes the ergotropy of the state, which has a long history in literature [40, 41, 42]. The corresponding function $\mathbb{H}_{GL}(\rho_{\bf A})$ quantifies the ergotropic gap – the difference between the extractable work from a state $\rho_{\bf A}$ under global and local unitary operations, respectively. Interestingly, this function and its variants can capture different forms of nonclassical correlations, particularly entanglement in bipartite states and genuine entanglement in multipartite quantum states [43, 44, 45, 46, 47].
Figure 2: Nonlocal energy locking and superadditivity phenomena. Consider the one parameter family of states $\rho^{\alpha}_{\bf A}$. Black dotted curve represents quantum discord $\mathcal{Q}(\rho^{\alpha}_{\bf A})$ [48]. Fixing $\epsilon_{2}=2$, the ergodiscord $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho^{\alpha}_{\bf A})$ is plotted for different values of $\epsilon_{1}$. For $\epsilon_{1}=6$, the value of ergodiscord exceeds unity for certain range of the parameter $\alpha$. Notably, the corresponding states lock more energy than the maximally entangled state |$\ket{\phi^{\max}}$. These states when compared with maximally entangled state, depicts the feature as stated in Theorem 4.
@@
While studying the quantity $\mathbb{H}_{G\mathcal{N}^{0}_{C}}$, which is one of our core interests in the present work, we first need to identify the free classical channels $\mathcal{N}^{0}_{C}$. As it can be argued a classical channel $N\in\mathcal{N}_{C}$ communicating $A_{j}$ sub-part to the $j^{\star}$ party will be considered to be free if $\operatorname{Tr}(h_{j}N(\rho_{j}))\leq\operatorname{Tr}(h_{j}\rho_{j})\leavevmode\nobreak\ \&\leavevmode\nobreak\ \Delta h^{U_{j}}_{j}N(\rho_{j})\leq\Delta h^{U_{j}}_{j}(\rho_{j}),\leavevmode\nobreak\ \forall\leavevmode\nobreak\ \rho_{j}\in\mathcal{D}(\mathbb{C}^{d_{j}})$. A more elaborate justification of this demand as well as a complete characterization of free classical channels $\mathcal{N}^{0}_{C}$ are provided in Appendix IV. With this we now present a general result for bipartite systems (proof is deferred to Appendix IV).
@@
The function $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$, termed ergodiscord, faithfully captures nonclassical correlation from $A_{1}\to A_{2}$ in any bipartite state $\rho_{\bf A}$, whenever $h_{\bf A}$ has non degenerate spectrum.
Notably unlike Theorem 1, the function $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ faithfully captures signature of nonclassical correlation in any bipartite state without assuming any further constraint on the state. Theorem 2 has quite interesting implications. A nonzero value of $\mathbb{H}_{G\mathcal{N}^{0}_{C}}(\rho_{\bf A})$ implies that the extractable work under global operations is strictly greater than the work extractable under local operations assisted with classical communication from $A_{1}$ to $A_{2}$. In other words, work gets locked in the nonclassical correlation of the state. This phenomenon we call the nonlocal energy locking in nonclassical correlation. The motivation of this name arises from the pioneering ‘nonlocality without entanglement’ phenomenon, where distinguishability of a set of multipartite product states under local operation and classical communication is suboptimal than the global distinguishability [49, 50] (see also [51, 52, 53, 54]), and the works [55, 56, 57] that report several locking phenomena utilizing quantum resources). The following crucial property of the quantity discussed in Theorem 2 is worth noting.